OpenStudy (anonymous):

How do you solve 9x squared-6x-2 =0?

OpenStudy (anonymous):

≈0.910683602523 ≈-0.2440169358563

OpenStudy (anonymous):

wait I don't get it... how do you solve it?

OpenStudy (anonymous):

\[x=(\sqrt{3}+a)/3 ; \approx.0.91\] \[x=(1-\sqrt{3})/3;\approx -0.2\]

OpenStudy (anonymous):

I don't think that is right, because that not one of the options.

OpenStudy (anonymous):

wait a minute I'm adjusting the response

OpenStudy (anonymous):

quadratic equation formula method \[9x ^{2} +(-6)x-2=0\] ; \[ x= (-(-6+\sqrt{(-6)^{2}-4*9(-2}))/2*9)\] ; \[x=(-(-6)+\sqrt{108})/2*9\] ; x=(-(-6)+\sqrt{36+(-36)(-2)})/2*9\] ; \[x=(-(-6)+6\sqrt{3}) /2*9 ; x=(6-6\sqrt{3})/18 ; x=1-\sqrt{3}/3\] ;\[x=\sqrt{3}+1/3; x=(1-\sqrt{3})/3 \]

OpenStudy (anonymous):

quadratic equation formula \[x=(-b+/-\sqrt{b ^{2}-4ac)}/2a\]